# AudioBench: A Universal Benchmark for Audio Large Language Models

### v1

@@ 3.1 Main Concept
Category Dataset Name # of Samples Total Hours Avg Len (s) Min Len (s) Max Len (s) Audio Understanding ASR LibriSpeech-Clean 2,617 5.40 7.43 1.28 34.96 LibriSpeech-Other 2,935 5.34 6.55 1.47 34.51 CommonVoice-15 16,348 26.95 5.93 1.34 105.67 PeoplesSpeech 32,603 59.20 6.54 1.00 99.91 GigaSpeech 18,650 35.09 6.77 1.00 22.03 Tedlium3 1,142 2.61 8.24 1.07 32.55 Tedlium3-Longform 11 2.61 856.51 340.96 1,596.73 Earning-21 44 39.26 3,212.41 1,097.10 5,740.61 Earning-22 125 119.88 3,452.71 874.73 7,407.05 SQA CN-College-Listen 2,271 13.3 21.09 5.76 137.82 SLUE-P2-SQA5 408 4.5 39.85 13.00 40.0 DREAM-TTS 1,913 18.1 34.14 3.14 261.85 Public-SG-SpeechQA 688 7.6 39.86 15.78 95.71 SI OpenHermes-Audio 100 0.16 5.95 2.04 15.77 ALPACA-Audio 100 0.12 4.32 1.80 8.85 Audio Scene Understanding AQA Clotho-AQA 2,251 14.1 22.59 15.03 29.97 WavCaps-QA 304 0.87 10.28 1.0 30.62 AudioCaps-QA 313 0.86 9.86 3.27 10.00 AC WavCaps 1,730 4.9 10.22 1.00 30.97 AudioCaps 4,400 12.1 9.86 1.74 10.0 Voice Understanding ER IEMOCAP-Emotion 1,002 1.3 4.51 0.75 24.12 MELD-Sentiment 2,610 2.4 3.35 0.13 304.96 MELD-Emotion 2,610 2.4 3.35 0.13 304.96 AR VoxCeleb1-Accent 4,874 11.2 8.27 3.96 69.04 GR VoxCeleb1-Gender 4,874 11.2 8.27 3.96 69.04 IEMOCAP-Gender 1,002 1.26 4.55 0.69 26.77
@@ 3.3.1 Speech Understanding
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Audio Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.70B 1.95 2.15 2.51 2.96 SLUE-P2-SQA5 3.95 3.84 4.16 4.19 DREAM-TTS 1.79 1.91 2.50 2.82 Public-SG-SpeechQA 2.77 2.97 2.98 3.25 OpenHermes-Audio 1.05 1.14 1.07 2.82 ALPACA-Audio 0.85 1.12 1.15 3.37 Audio Scene Understanding Clotho-AQA M.J.70B 2.75 2.88 2.16 1.40 WavCaps-QA 2.31 1.97 1.34 0.83 AudioCaps-QA 2.39 2.41 1.34 0.77 WavCaps 1.25 1.41 0.49 0.23 AudioCaps 1.83 2.11 0.50 0.24 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.70B 2.33 1.60 2.43 1.85 MELD-Sentiment 1.82 1.72 1.65 2.26 MELD-Emotion 1.60 1.46 1.45 1.88 VoxCeleb1-Accent 1.50 2.25 1.87 1.19 VoxCeleb1-Gender 3.98 3.59 2.96 2.87 IEMOCAP-Gender 2.02 2.55 2.01 2.51
@@ Appendix A Appendix
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Audio Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.8B 1.57 1.63 1.95 2.32 SLUE-P2-SQA5 3.09 3.03 3.08 3.26 DREAM-TTS 1.46 1.35 1.86 2.23 Public-SG-SpeechQA 2.18 2.42 2.18 2.51 OpenHermes-Audio 2.15 2.12 1.72 2.34 ALPACA-Audio 1.72 2.19 1.71 2.39 Audio Scene Understanding Clotho-AQA M.J.8B 2.08 2.26 1.76 1.28 WavCaps-QA 1.90 1.58 1.03 0.95 AudioCaps-QA 1.94 1.93 1.06 1.05 WavCaps 1.18 1.16 0.51 0.65 AudioCaps 1.47 1.39 0.49 0.56 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.8B 1.59 1.00 1.71 1.44 MELD-Sentiment 1.51 1.21 1.30 1.93 MELD-Emotion 1.16 1.02 1.04 1.47 VoxCeleb1-Accent 1.52 1.59 1.38 0.95 VoxCeleb1-Gender 2.58 2.74 2.41 1.67 IEMOCAP-Gender 1.35 1.95 1.54 1.56

### v2

@@ 3.1 Main Concept
Category Dataset Name # of Samples Total Hours Avg Len (s) Min Len (s) Max Len (s) Speech Understanding ASR LibriSpeech-Clean 2,617 5.40 7.43 1.28 34.96 LibriSpeech-Other 2,935 5.34 6.55 1.47 34.51 CommonVoice-15 16,348 26.95 5.93 1.34 105.67 PeoplesSpeech 32,603 59.20 6.54 1.00 99.91 GigaSpeech 18,650 35.09 6.77 1.00 22.03 Tedlium3 1,142 2.61 8.24 1.07 32.55 Tedlium3-Longform 11 2.61 856.51 340.96 1,596.73 Earning-21 44 39.26 3,212.41 1,097.10 5,740.61 Earning-22 125 119.88 3,452.71 874.73 7,407.05 SQA CN-College-Listen 2,271 13.3 21.09 5.76 137.82 SLUE-P2-SQA5 408 4.5 39.85 13.00 40.0 DREAM-TTS 1,913 18.1 34.14 3.14 261.85 Public-SG-SpeechQA 688 7.6 39.86 15.78 95.71 SI OpenHermes-Audio 100 0.16 5.95 2.04 15.77 ALPACA-Audio 100 0.12 4.32 1.80 8.85 Audio Scene Understanding AQA Clotho-AQA 2,251 14.1 22.59 15.03 29.97 WavCaps-QA 304 0.87 10.28 1.0 30.62 AudioCaps-QA 313 0.86 9.86 3.27 10.00 AC WavCaps 1,730 4.9 10.22 1.00 30.97 AudioCaps 4,400 12.1 9.86 1.74 10.0 Voice Understanding ER IEMOCAP-Emotion 1,002 1.3 4.51 0.75 24.12 MELD-Sentiment 2,610 2.4 3.35 0.13 304.96 MELD-Emotion 2,610 2.4 3.35 0.13 304.96 AR VoxCeleb1-Accent 4,874 11.2 8.27 3.96 69.04 GR VoxCeleb1-Gender 4,874 11.2 8.27 3.96 69.04 IEMOCAP-Gender 1,002 1.26 4.55 0.69 26.77
@@ 3.3.1 Speech Understanding
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Speech Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.70B 1.95 2.15 2.51 2.96 SLUE-P2-SQA5 3.95 3.84 4.16 4.19 DREAM-TTS 1.79 1.91 2.50 2.82 Public-SG-SpeechQA 2.77 2.97 2.98 3.25 OpenHermes-Audio 1.05 1.14 1.07 2.82 ALPACA-Audio 0.85 1.12 1.15 3.37 Audio Scene Understanding Clotho-AQA M.J.70B 2.75 2.88 2.16 1.40 WavCaps-QA 2.31 1.97 1.34 0.83 AudioCaps-QA 2.39 2.41 1.34 0.77 WavCaps 1.25 1.41 0.49 0.23 AudioCaps 1.83 2.11 0.50 0.24 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.70B 2.33 1.60 2.43 1.85 MELD-Sentiment 1.82 1.72 1.65 2.26 MELD-Emotion 1.60 1.46 1.45 1.88 VoxCeleb1-Accent 1.50 2.25 1.87 1.19 VoxCeleb1-Gender 3.98 3.59 2.96 2.87 IEMOCAP-Gender 2.02 2.55 2.01 2.51
@@ Appendix A Appendix
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Speech Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.8B 1.57 1.63 1.95 2.32 SLUE-P2-SQA5 3.09 3.03 3.08 3.26 DREAM-TTS 1.46 1.35 1.86 2.23 Public-SG-SpeechQA 2.18 2.42 2.18 2.51 OpenHermes-Audio 2.15 2.12 1.72 2.34 ALPACA-Audio 1.72 2.19 1.71 2.39 Audio Scene Understanding Clotho-AQA M.J.8B 2.08 2.26 1.76 1.28 WavCaps-QA 1.90 1.58 1.03 0.95 AudioCaps-QA 1.94 1.93 1.06 1.05 WavCaps 1.18 1.16 0.51 0.65 AudioCaps 1.47 1.39 0.49 0.56 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.8B 1.59 1.00 1.71 1.44 MELD-Sentiment 1.51 1.21 1.30 1.93 MELD-Emotion 1.16 1.02 1.04 1.47 VoxCeleb1-Accent 1.52 1.59 1.38 0.95 VoxCeleb1-Gender 2.58 2.74 2.41 1.67 IEMOCAP-Gender 1.35 1.95 1.54 1.56